# Water-based Quantum Dots Liquid Scintillator for Particle Physics

### v1

@@ Abstract
Liquid scintillators are typically composed from organic compounds dissolved in organic solvents. However, usage of such material is often restricted due to fire safety and environmental reasons. Because of this, R&D of water-based liquid scintillators is of extreme relevance; yet, no such scintillators have been made commercially available as yet. Here, we investigate an alternative, water-based quantum dots liquid scintillator. Pre-determined and controllable optical properties of the quantum dots, as well as the existence of large libraries of established protocols for their dispersion in aqueous solutions, make them an attractive option for nuclear and particle physics applications. We characterize the optical properties of water-based quantum dots liquid scintillator and find that most of its optical properties are preserved upon quantum dots’ phase transfer into water, through the addition of an oleic acid hydrophilic layer. Using the developed scintillator, the time and charge responses from atmospheric muons are measured, highlighting the practical viability of water-based quantum dots liquid scintillators for nuclear and particle physics, special interest on neutrino physics.
@@ 1 Introduction
Here we are investigating the possibility of utlizing QDs for particle physics applications [13, 14, 15]. In particular, we are interested in the possibility of WbQDs liquid scintillator in nuclear fission reactor neutrino detection. Water is a safe material to use near the fission reactor core, and such detector can be a new type of near-field fission reactor monitoring detector [16, 17, 18, 19, 20]. Typical compositions of the QDs include elements such as cadmium, selenium, and tellurium, all of which have high neutron capture cross-sections. This opens a possibility to tag neutrons, and WbQDs to act as conventional liquid scintillators with neutron detection ability without doping neutron capturing materials, in contrast to the proposed gadolinium-doped water-based liquid scintillator [20].
@@
Regarding the emission properties of QDs, the vast majority of photodetectors used for radiation detection have maximum quantum efficiencies (QE) in the blue end of the spectrum; therefore, the selection of QDs’ composition and size needs to be such as to yield emission in this spectral range. QDs with Cd-based compositions and small core sizes seem appropriate for proof-of-principle demonstrations. Furthermore, and although not explored in this work, cadmium has a high-neutron capture cross-section and Cd-based cores of our QDs open a possibility to employ them as neutron capture materials, allowing tagging of neutrino interactions through the inverse beta-decay processes [22]. Regardless of the application, the concentration of QDs in the scintillating liquid needs to be high enough to facilitate sufficient muon interactions or neutron capture events for experimentation, whilst low enough to maintain the optical transparency of the liquid. As such, there will be an optimum range for QD concentration in WbQD scintillating liquids.
Finally, for application purposes, QDs must remain colloidally and functionally stable for relatively long periods of time. Whilst it is known that phase-transferred QDs might gradually loose their stability over time and that routes to mitigate this do exist [23, 24, 25, 26], little is known about the effect of long-term exposure to radiation on the QDs’ stability and functionality. Although radiation responses of QDs were studied in the past [9, 10], these were for low energy particles and not MIPs such as cosmic rays. As such, measuring MIP responses and therefore demonstrating the scintillating property of WbQDs liquid scintillator - the subject the work presented here - is the first step towards the utilization of WbQDs as particle scintillators.
@@ 2.1 Quantum dots selected
To meet the required characteristics discussed in the previous section, cadmium sulfide (CdS) core zinc sulfide (ZnS) shell type coating quantum dots were selected for the experiments. There were numerous reasons behind this selection. Firstly, the CdS composition of the QD core facilitates shorter wavelengths of QD fluorescence, compared to e.g. the more popular CdSe or CdTe QDs. Secondly, blue-emitting CdS QDs of relatively high PLQY ( $\sim$50%) were available for purchase commercially  [27], although these QDs could not be dispersed in water directly, requiring further surface modification to transfer them to aqueous environments (as detailed in sec. 2.2). Thirdly, the manufacturer stated the shelf-life of these QDs as being three years, which was sufficient for our purposes due to the short time span of the experiment. And finally, the cadmium isotope ${}^{113}$d has a natural abundance of $12\%$, and a relatively high neutron capture cross-section of $2.1\times 10^{4}$ barn [28], with a Q-value $9.04$ MeV.
@@ 2.2 Surfactant encapsulation method
Three samples were prepared in this manner, one in each of the years of 2021, 2022 and 2023, and labelled accordingly. Figure 1 shows the three samples under natural and UV illumination (recorded in August of 2023). The blue color of solutions under UV illumination (in panel (b)) is indicative of QD presence in the dispersions and therefore of the successful completion of the phase transfer procedure on each occasion. It is clear that in all three cases the concentration of QDs is relatively high; yet, and as is evident from panel (a) of Figure 1, all three samples have a high degree of transparency, as per requirements for successful neutrino experiments. In the following result section, we present a representative set of data from the 2023 sample; the properties of 2021 and 2022 samples will be discussed in detail in the section 3.3, where sample aging is considered.
@@ 3.1 Optical properties of CdS/ZnS QDs
The extinction and photoluminescence (PL) spectra of QD dispersions were recorded using 2600i UV-vis spectrometer from Shimadzu [32] and Cary Eclipse Fluorescence Spectrophotometer [33] respectively. In our interpretation of the extinction data, we attributed any strong changes in the extinction of the sample to the absorption by the QDs, since they have relatively small scattering cross-sections [34]. However, below the band gap of the QDs, scattering contributions can and do become dominant.
Figure 2: Optical properties of CdS/ZnS QDs. Panels (a) and (b) show absorbance (solid lines, left scales in absorbance unit or a.u.) and photoluminescence (dashed lines, right scales in arbitrary normalization) of (a) toluene- and (b) water-dispersed QDs. Panel (c) compares the normalized photoluminescence spectra of the toluene- and water-dispersed QDs samples. Panel (d) shows the experimental data (black scatter points) for the water-dispersed QD sample and the Gaussian fit to the data, that was used to determine the position and width of the first excitonic peak of the QDs. The $R^{2}$ value for the fit was 0.99983 and the corresponding $\chi^{2}$ value was $1.2\times 10^{-8}$.
As expected, and as shown in Figs. 2(a) and (b), the absorbance of QDs dispersions in toluene and in water remains negligible at low energies (high wavelengths), but increases rapidly as the energy of the excitation becomes greater than the bandgap of the QDs (at $\sim$470 nm for both). No significant changes to the absorption spectra could be observed for the WbQD liquid as compared to the toluene-dispersed QDs, indicating that no significant chemical or physical changes to the QD core ocurred as a result of the phase transfer process. However, comparison of the extinction spectra of the two samples at wavelengths above the QDs’ band edge, it is clear that the phase transfer of QDs to an aqueous environment resulted in an increased Rayleigh’s scattering by the QDs - this appears as an elevated but gradually decreasing background at wavelengths larger than 470 nm. This is a typical observation, owing to the size dependency of the Rayleigh scattering process and simply indicates an increased overall size of the QDs upon addition of the surfactant coating that is needed to achieve water-dispersibility [35, 36].
@@
Figure 3: Determination of PLQY of WbQD liquid. Panels (a) and (b) show photoluminescence spectra of (a) Atto390 and (b) WbQD samples of seven different concentrations. Absorbance of each sample at the excitation wavelength ($\lambda_{exc}=$380 nm) is stated in the legend of each panel. (c) Extinction spectra of Atto390 (black line) and WbQD samples (red line). The excitation wavelength used for acquisition of photoluminescence spectra, shown in panels (a) and (b), is indicted by the gray dashed line. (d) Integrated photoluminescence intensity of Atto390 (black data) and CdS QDs (red data) as a function of sample absorbance at 380 nm. The experimental data are shown as scatter points. The lines are linear fits to the data; the $R^{2}$ value of each fit and the slope obtained from it are indicated next to the lines in respective colours. The ratio of the slopes is proportional to the ratio of the quantum yields of the two samples.
@@
Sample
Slope, $K$
Intercept, $I_{0}$
$R^{2}$ value
Atto390 dye
$(4144\pm 124)\times 10^{3}$
$(-5.25\pm 5.98)\times 10^{3}$
0.9955
2023 WbQDs
$(437\pm 9)\times 10^{3}$
$(-1.30\pm 0.44)\times 10^{3}$
0.9978
@@
$A_{0}$
$A_{max}$
$\lambda_{c}$
$w$
$FWHM$
Value
0.00947
0.0229
448.8
11.6
27.3
Error
0.00006
0.0002
0.3
0.2
0.4
@@ 3.2 Physical properties of CdS/ZnS QDs
Figure 4(a) shows the normalized correlation curves of individual measurements as well as their average for CdS QDs dispersed in toluene and in water, after phase transfer. From this data, it is clear that correlation in scattered signals persists for longer lag times for WbQDs, which is indicative of their increased hydrodynamic size as compared to QDs dispersed in toluene. This was confirmed by the cumulant analysis, which yielded mean QD sizes of $25\pm 9$ nm and $71.6\pm 0.3$ nm pre- and post-phase transfer, with the population standard deviations of $11$ nm and $26$ nm respectively. The size distribution data, shown in Figure 4(b) further confirms these results - the hydrodynamic size of WbQDs increases strongly compared to their predecessors dispersed in toluene. It is important to note that the observed large increase in the size of the QDs (approx. 3 fold) is due to both their increased physical size, owing to the additional OA layer around each QD, but also due to the emergence of charges on their surfaces which interact with the surrounding polar water molecules. It is known that the electrostatic force between charged particles and the polar water molecules contributes to the friction experienced by the particle when diffusing, resulting in larger effective hydrodynamic sizes [47, 48]. We confirmed the existence of surface charges on the phase-transfered QDs by measuring their zeta potential using DLS. Figure 4(c) shows the results of these measurements for WbQD sample. Analysis of these results (performed automatically by the Zetasizer software) yielded a mean zeta-potential value of
@@
Finally, the physical size of the QDs’ semiconductor core was determined using Transmission Electron Microscopy (TEM). A representative TEM image obtain from the measurements in shown in Fig. 4, which was analysed to obtain the mean QD diameter of $5.0\pm 1.6$ nm. This value is comparable to that obtained from Eq. 3.5, and any discrepancy between the two can be accounted for by the limited numbers of QDs that could be imaged by the TEM technique.
@@ 3.3 Effects of sample ageing
Figure 5: Effects of sample ageing on the spectral properties of phase-transferred CdSe QDs. Panels (a) and (b) show absorbance (black lines, left scales) and fluorescence (red lines, right scales) spectra of CdS QDs phase-transferred in July of (a) 2021 and (b) 2022. The top graphs in each panel show the optical spectra recorded for CdS samples immediately after phase-transfer process, and the bottom graphs show the same set of data measured in July of 2023. The gray dashed line in each panel indicates the peak wavelength of the freshly-transferred QDs’ emission.
@@
Sample
PLQY (%)
$D_{h}$ (nm)
$\zeta$ (mV)
2021 WbQD
as prepared
$15.6\pm 1.2$
$52.2\pm 0.3$
not available
in 2023
$8.8\pm 0.8$
$48.7\pm 0.1$
$-13.8\pm 0.8$
2022 WbQD
as prepared
$8.8\pm 1.5$
$60\pm 1$
$-24\pm 2$
in 2023
$14.5\pm 0.7$
$55.7\pm 0.3$
$-47\pm 2$
2023 WbQD
as prepared
$9.5\pm 0.5$
$71.6\pm 0.3$
$-24.6\pm 0.7$
@@ 4.1 Setup
Figure 6: (a) Cosmic ray test diagram. 2 cosmic ray taggers are piled to define a cosmic ray trajectory. A bottle of CdS QDs in water solution is placed in a dark box where 3-inch PMT is located in front of it. Coincidence signals from the cosmic ray taggers are used to trigger the digitizer to record PMT pulses. (b) A view of the inside of the dark box. Note, cosmic ray tagger (the blue box) locates outside of the dark box unlike this picture.
@@ 4.2 Simulation
The geometry created is a simplified replica of the physical setup taking into consideration the distance, size, and shape of the objects. Fig. 8 left shows an event display of our simulation in GEANT4. The PMT hemispherical glass surface is set as the sensitive detector. The WbQD material is treated as a liquid scintillator with close emission spectrum, located inside of a cylindrical glass. Utilising the necessary GEANT4 libraries such as electromagnetic processes, the physics processes can be constructed. Muons are generated randomly and they penetrate the WbQDs volume and produce scintillation photons that are detected by the PMT. The trajectories of muons are calculated by the solid angle defined by 2 cosmic ray taggers, and for simplicity all muons are 4 GeV in this simulation. In the simulation, we do not simulate PMT responses, instead, all photons which hit the surface of the PMT are registered with constant $15\%$ efficiency and the gain is $1.0\times 10^{7}$ at $-1100$ V according to our calibration.
Fig. 8 right shows data-MC comparison of their charge distributions above the pedestal. Here, the number of scintillation photons in the simulation is determined by adjusting the two distributions to match. A good agreement between experimental and simulated data is achieved around $4000$ photons/MeV, although with some deviations for low charge region. This is likely due to the contamination of background events in the data, and further investigations are required to confirm this. The simulation adopts a simple geometry, and objects in the dark box, except those shown in Fig. 8, left, are ignored. However, we believe they do not affect the time and charge distributions much, as major signals are direct light travelling from the WbQD scintillator to the front of the PMT. Importantly, this is the first demonstration of the water-based quantum dots liquid scintillator with radiation detection; furthermore, the number of photons recorded here is promising for applications in particle and nuclear physics.
Figure 8: Simulation of the experimental setup in Geant4. (a) Event display of the simulation. Cosmic ray distributions are based on the configuration of 2 cosmic ray taggers and example muon trajectories are shown in green. (b) Data-simulation charge distribution comparison. Data pedestal is cut, and number of scintillation photons per MeV is adjusted so that the distributions agree.

### v2

@@ Abstract
Liquid scintillators are typically composed from organic compounds dissolved in organic solvents. However, usage of such material is often restricted due to fire safety and environmental reasons. Because of this, R&D of water-based liquid scintillators is of extreme relevance; yet, no such scintillators have been made commercially available as yet. Here, we investigate an alternative, water-based quantum dots liquid scintillator. Pre-determined and controllable optical properties of the quantum dots, as well as the existence of large libraries of established protocols for their dispersion in aqueous solutions, make them an attractive option for nuclear and particle physics applications. We characterize the optical properties of water-based quantum dots liquid scintillator and find that most of its optical properties are preserved upon quantum dots’ phase transfer into water, through the addition of an oleic acid hydrophilic layer. Using the developed scintillator, the time and charge responses from atmospheric muons are measured, highlighting the practical viability of water-based quantum dots liquid scintillators for nuclear and particle physics, with special interest in neutrino physics.
@@ 1 Introduction
Here we are investigating the possibility of utlizing QDs for particle physics applications [13, 14, 15]. In particular, we are interested in the possibility of WbQDs liquid scintillator in nuclear fission reactor neutrino detection. Water is a safe material to use near the fission reactor core, and such a detector can be a new type of near-field monitor for fission reactors [16, 17, 18, 19, 20, 21, 22]. Typical compositions of the QDs include elements such as cadmium, selenium, and tellurium, all of which have high neutron capture cross-sections. This opens a possibility to tag neutrons, and WbQDs to act as conventional liquid scintillators with neutron detection ability without doping neutron capturing materials, in contrast to the proposed gadolinium-doped water-based liquid scintillator [20].
@@
Regarding the emission properties of QDs, the vast majority of photodetectors used for radiation detection have maximum quantum efficiencies (QE) in the blue end of the spectrum; therefore, the selection of QDs’ composition and size needs to be such as to yield emission in this spectral range. QDs with Cd-based compositions and small core sizes seem appropriate for proof-of-principle demonstrations. Furthermore, and although not explored in this work, cadmium has a high neutron capture cross-section and Cd-based cores of our QDs open a possibility to employ them as neutron capture materials, allowing tagging of neutrino interactions through the inverse beta-decay processes [24]. Regardless of the application, the concentration of QDs in the scintillating liquid needs to be high enough to facilitate sufficient muon interactions or neutron capture events for experimentation, whilst low enough to maintain the optical transparency of the liquid. As such, there will be an optimum range for QD concentration in WbQD scintillating liquids.
Finally, for application purposes, QDs must remain colloidally and functionally stable for relatively long periods of time. Whilst it is known that phase-transferred QDs might gradually lose their stability over time and that routes to mitigate this do exist [25, 26, 27, 28], little is known about the effect of long-term exposure to radiation on the QDs’ stability and functionality. Although radiation responses of QDs were studied in the past [9, 10], these were for low energy particles and not MIPs such as cosmic rays. As such, measuring MIP responses and therefore demonstrating the scintillating property of WbQDs liquid scintillator - the subject the work presented here - is the first step towards the utilization of WbQDs as particle scintillators.
@@ 2.1 Quantum dots selected
To meet the required characteristics discussed in the previous section, cadmium sulfide (CdS) core zinc sulfide (ZnS) shell type coating quantum dots were selected for the experiments. There were numerous reasons behind this selection. Firstly, the CdS composition of the QD core facilitates shorter wavelengths of QD fluorescence, compared to e.g. the more popular CdSe or CdTe QDs. Secondly, blue-emitting CdS QDs of relatively high PLQY ( $\sim$50%) were available for purchase commercially  [29], although these QDs could not be dispersed in water directly, requiring further surface modification to transfer them to aqueous environments (as detailed in sec. 2.2). Thirdly, the manufacturer stated the shelf-life of these QDs as being three years, which was sufficient for our purposes due to the short time span of the experiment. And finally, the cadmium isotope 113Cd has a natural abundance of $12\%$, and a relatively high neutron capture cross-section of $2.1\times 10^{4}$ barn [30], with a Q-value $9.04$ MeV.
@@ 2.2 Surfactant encapsulation method
Three samples were prepared in this manner, one in each of the years of 2021, 2022 and 2023, and labelled accordingly. Figure 1 shows the three samples under natural and UV illumination (recorded in August of 2023). The blue color of solutions under UV illumination (in panel (b)) is indicative of QD presence in the dispersions and therefore of the successful completion of the phase transfer procedure on each occasion. It is clear that in all three cases the concentration of QDs is relatively high; yet, and as is evident from panel (a) of Figure 1, all three samples have a high degree of transparency, as per requirements for successful neutrino experiments. In the following result section, we present a representative set of data from the 2023 sample; the properties of 2021 and 2022 samples will be discussed in detail in section 3.3, where sample aging is considered.
@@ 3.1 Optical properties of CdS/ZnS QDs
The extinction and photoluminescence (PL) spectra of QD dispersions were recorded using a 2600i UV-vis spectrometer from Shimadzu [34] and Cary Eclipse Fluorescence Spectrophotometer [35] respectively. In our interpretation of the extinction data, we attributed any strong changes in the extinction of the sample to the absorption by the QDs, since they have relatively small scattering cross-sections [36]. However, below the band gap of the QDs, scattering contributions can and do become dominant.
Figure 2: Optical properties of CdS/ZnS QDs. Panels (a) and (b) show absorbance (solid lines, left scales in absorbance unit or a.u.) and photoluminescence (dashed lines, right scales in arbitrary normalization) of (a) toluene- and (b) water-dispersed QDs. Panel (c) compares the normalized photoluminescence spectra of the toluene- and water-dispersed QDs samples. Panel (d) shows the experimental data (black scatter points) for the water-dispersed QD sample and the Gaussian fit to the data that was used to determine the position and width of the first excitonic peak of the QDs. The $R^{2}$ value for the fit was 0.99983 and the corresponding $\chi^{2}$ value was $1.2\times 10^{-8}$.
As expected, and as shown in Figs. 2(a) and (b), the absorbance of QDs dispersions in toluene and in water remains negligible at low energies (high wavelengths), but increases rapidly as the energy of the excitation becomes greater than the bandgap of the QDs (at $\sim$470 nm for both). No significant changes to the absorption spectra could be observed for the WbQD liquid as compared to the toluene-dispersed QDs, indicating that no significant chemical or physical changes to the QD core occurred as a result of the phase transfer process. However, comparison of the extinction spectra of the two samples at wavelengths above the QDs’ band edge, it is clear that the phase transfer of QDs to an aqueous environment resulted in an increased Rayleigh’s scattering by the QDs - this appears as an elevated but gradually decreasing background at wavelengths larger than 470 nm. This is a typical observation, owing to the size dependency of the Rayleigh scattering process and simply indicates an increased overall size of the QDs upon addition of the surfactant coating that is needed to achieve water-dispersibility [37, 38].
@@
Figure 3: Determination of PLQY of WbQD liquid. Panels (a) and (b) show photoluminescence spectra of (a) Atto390 and (b) WbQD samples of seven different concentrations. Absorbance of each sample at the excitation wavelength ($\lambda_{exc}=$380 nm) is stated in the legend of each panel. (c) Extinction spectra of Atto390 (black line) and WbQD samples (red line). The excitation wavelength used for acquisition of photoluminescence spectra, shown in panels (a) and (b), is indicated by the gray dashed line. (d) Integrated photoluminescence intensity of Atto390 (black data) and CdS QDs (red data) as a function of sample absorbance at 380 nm. The experimental data are shown as scatter points. The lines are linear fits to the data; the $R^{2}$ value of each fit and the slope obtained from it are indicated next to the lines in respective colours. The ratio of the slopes is proportional to the ratio of the quantum yields of the two samples.
@@
Sample Slope, $K$ Intercept, $I_{0}$ $R^{2}$ value
Atto390 dye $(4144\pm 124)\times 10^{3}$ $(-5.25\pm 5.98)\times 10^{3}$ 0.9955
2023 WbQDs $(437\pm 9)\times 10^{3}$ $(-1.30\pm 0.44)\times 10^{3}$ 0.9978
@@
$A_{0}$ $A_{max}$ $\lambda_{c}$ $w$ $FWHM$
Value 0.00947 0.0229 448.8 11.6 27.3
Error 0.00006 0.0002 0.3 0.2 0.4
@@ 3.2 Physical properties of CdS/ZnS QDs
Figure 4(a) shows the normalized correlation curves of individual measurements as well as their average for CdS QDs dispersed in toluene and in water, after phase transfer. From this data, it is clear that correlation in scattered signals persists for longer lag times for WbQDs, which is indicative of their increased hydrodynamic size as compared to QDs dispersed in toluene. This was confirmed by the cumulant analysis, which yielded mean QD sizes of $25\pm 9$ nm and $71.6\pm 0.3$ nm pre- and post-phase transfer, with the population standard deviations of $11$ nm and $26$ nm respectively. The size distribution data, shown in Figure 4(b) further confirms these results - the hydrodynamic size of WbQDs increases strongly compared to their predecessors dispersed in toluene. It is important to note that the observed large increase in the size of the QDs (approx. 3 fold) is due to both their increased physical size, owing to the additional OA layer around each QD, but also due to the emergence of charges on their surfaces which interact with the surrounding polar water molecules. It is known that the electrostatic force between charged particles and the polar water molecules contributes to the friction experienced by the particle when diffusing, resulting in larger effective hydrodynamic sizes [49, 50]. We confirmed the existence of surface charges on the phase-transferred QDs by measuring their zeta potential using DLS. Figure 4(c) shows the results of these measurements for the WbQD sample. Analysis of these results (performed automatically by the Zetasizer software) yielded a mean zeta-potential value of
@@
Finally, the physical size of the QDs’ semiconductor core was determined using Transmission Electron Microscopy (TEM). A representative TEM image obtained from the measurements is shown in Fig. 4, which was analysed to obtain the mean QD diameter of $5.0\pm 1.6$ nm. This value is comparable to that obtained from Eq. 3.5, and any discrepancy between the two can be accounted for by the limited numbers of QDs that could be imaged by the TEM technique.
@@ 3.3 Effects of sample ageing
Figure 5: Effects of sample ageing on the spectral properties of phase-transferred CdSe QDs. Panels (a) and (b) show absorbance (black lines, left scales) and fluorescence (red lines, right scales) spectra of CdS QDs phase-transferred in July of (a) 2021 and (b) 2022. The top graphs in each panel show the optical spectra recorded for CdS samples immediately after the phase-transfer process, and the bottom graphs show the same set of data measured in July of 2023. The gray dashed line in each panel indicates the peak wavelength of the freshly-transferred QDs’ emission.
@@
Sample PLQY (%) $D_{h}$ (nm) $\zeta$ (mV)
2021 WbQD as prepared $15.6\pm 1.2$ $52.2\pm 0.3$ not available
in 2023 $8.8\pm 0.8$ $48.7\pm 0.1$ $-13.8\pm 0.8$
2022 WbQD as prepared $8.8\pm 1.5$ $60\pm 1$ $-24\pm 2$
in 2023 $14.5\pm 0.7$ $55.7\pm 0.3$ $-47\pm 2$
2023 WbQD as prepared $9.5\pm 0.5$ $71.6\pm 0.3$ $-24.6\pm 0.7$
@@ 4.1 Setup
Figure 6: (a) Cosmic ray test diagram. 2 cosmic ray taggers are piled to define a cosmic ray trajectory. A bottle of CdS QDs in a water solution is placed in a dark box with a 3-inch PMT located in front of it. Coincidence signals from the cosmic ray taggers are used to trigger the digitizer to record PMT pulses. (b) A view of the inside of the dark box. Note, the cosmic ray tagger (the blue box) is located outside of the dark box unlike this picture.
@@ 4.2 Simulation
The geometry created is a simplified replica of the physical setup taking into consideration the distance, size, and shape of the objects. Fig. 8 left shows an event display of our simulation in GEANT4. The PMT hemispherical glass surface is set as the sensitive detector. The WbQD material is treated as a liquid scintillator with a similar emission spectrum, located inside of a cylindrical glass. Utilising the necessary GEANT4 libraries such as electromagnetic processes, the physics processes can be constructed. Muons are generated randomly and they penetrate the WbQDs volume and produce scintillation photons that are detected by the PMT. The trajectories of muons are calculated by the solid angle defined by 2 cosmic ray taggers, and for simplicity all muons are 4 GeV in this simulation. In the simulation, we do not simulate PMT responses, instead, all photons which hit the surface of the PMT are registered with constant $15\%$ efficiency and the gain is $1.0\times 10^{7}$ at $-1100$ V according to our calibration.
Fig. 8 right shows data-MC comparison of their charge distributions above the pedestal. Here, the number of scintillation photons in the simulation is determined by adjusting the two distributions to match. A good agreement between experimental and simulated data is achieved around $4000$ photons/MeV, although with some deviations in the low charge region. This is likely due to the contamination of background events in the data, and further investigations are required to confirm this. The simulation adopts a simple geometry, and objects in the dark box, except those shown in Fig. 8, left, are ignored. However, we believe they do not affect the time and charge distributions much, as major signals are direct light travelling from the WbQD scintillator to the front of the PMT. Importantly, this is the first demonstration of the water-based quantum dots liquid scintillator with radiation detection; furthermore, the number of photons recorded here is promising for applications in particle and nuclear physics.
Figure 8: Simulation of the experimental setup in Geant4. (a) Event display of the simulation. Cosmic ray distributions are based on the configuration of 2 cosmic ray taggers and example muon trajectories are shown in green. (b) Data-simulation charge distribution comparison. Data pedestal is cut, and the number of scintillation photons per MeV is adjusted so that the distributions agree.